OpenStudy (anonymous):

how would you graph this 3y+4x=12

OpenStudy (amistre64):

ax+by=c xint = c/a yint = c/b

OpenStudy (amistre64):

plot those 2 marks and draw a line over them

OpenStudy (anonymous):

draw y=4-4x/3

OpenStudy (anonymous):

THANK YOU FOR YOUR HELP

OpenStudy (anonymous):